OpenStudy (ujjwal):

The sum of an infinite GS is 15 and the sum of their square is 45, then the common ratio is: a)3/4 b)1/4 c)2/3 d)5/6

OpenStudy (experimentx):

solve these \[ 15 = \frac{a}{1 - r}\] \[ 45 = \frac{a^2}{1 - r^2}\]

OpenStudy (ujjwal):

Thank you..